A 1,200 kg dragster, starting from rest, reaches a maximum velocity of 140 m/s in 5 seconds. At the 5 second mark, the dragster deploys a parachute bringing it to a stop in 3.5 seconds. Calculate the acceleration of the dragster after the parachute is deployed.
1
Expert's answer
2019-12-12T06:18:57-0500

The  acceleration of the dragster after the parachute is deployed is


"a=\frac{0-v}{t}=-\frac{v}{t}"

"a=-\frac{140}{3.5}=-40\frac{m}{s^2}"